# Critical Mathematical Economics and Progressive Data Science

### v2

@@ Abstract
From our point of view, mathematics has been partly misused in mainstream economics to justify ‘unregulated markets’ before the financial crisis. We thus identify two key parts of CME, which leads to a natural structure of this article: The first focusses on an analysis and critique of mathematical models used in mainstream economics, like e.g. the Dynamic Stochastic General Equilibrium (DSGE) in Macroeconomics and the so-called “Sonnenschein-Mantel-Debreu”-Theorems.
The aim of the second part is to improve and extend heterodox models using ingredients from modern mathematics and computer science, a method with strong relation to Complexity Economics. We exemplify this idea by describing how methods from Non-Linear Dynamics have been used in what could be called “The Dynamical Systems approach to Post-Keynesian Macroeconomics”, and also discuss (Pseudo-) Goodwin cycles and possible Micro- and Mesofoundations.
We conclude by discussing in which areas a collaboration between mathematicians and heterodox economists could be most promising. The focus lies on the mathematical and model-theoretic foundations of controversies in economic policy, and we discuss both existing projects in such a direction as well as areas where new models for policy advice are most needed. In an outlook, we discuss the role of (ecological) data, and the need for what we call Progressive Data Science.
@@ 1 Introduction
Recently, there has been a “Workshop on Complexity and Economics” ([57]) at the Fields Institute, which was organized in collaboration with the “Young Scholars Initiative” (YSI) of INET ([55]) and especially the “YSI Complexity Economics Working Group” ([56]. The aim of this article is to describe some of the research is this direction - in order to see what mathematics and mathematicians can contribute to the development of a competitive alternative to the neoclassical paradigm in economics.
2 The Use of Mathematics in Standard Macroeconomics
@@ 3.3 A probabilistic Micro-/Meso-foundation of Macroeconomics
4 Conclusion
@@
From a Dynamical Systems point of view, other promising topics to look at include Hysteresis, e.g. with relation to Keynesian modelling of the labour market ([49]). In the future maybe even advanced topics like Spatio-Temporal Patterns, Turing instability or Delayed Feedback Control could be part of the mathematics used for new economic thinking - for a panorama of recent advances in dynamical systems theory and their interplay with a wide range of applications see e.g. ([58]).
5 Outlook I: Foundations of Controversies in Economic Policy
@@
But a didactic presentation of the foundations of controversies in economic policy would also be very helpful to illuminate debates among progressive forces - let us take the Euro-Crisis and the question of the common currency as one example. Many conferences were held concerning the Plan B, which was introduced first by Jean-Luc Mélenchon in September 2015 ([60]). This Plan B would be implemented in case the Plan A for a fundamental democratization of the European Union was not feasible. The Plan B proposes that countries of the EU should have the possibility to leave the Euro-Zone and to return to their own national currencies. Others claim that the focus on the currency question is connected to substantial problems. In their view, on the one hand, the risks of leaving the common currency are underestimated, while, on the other hand, the leeway that is to be gained is overestimated (see e.g. [61]) For a panorama of the different positions discussed in the German Left, see e.g. [62] - we propose to base these discussions on a common framework of mathematical models of the (political) economy of Europe, in order to illuminate the debates and maybe even to find new roads for reconciling different political positions. This might become of crucial importance when the parties in the spectrum of social democrats, greens and democratic socialists will try to form progressive governments in the future. And such a question is obvoiusly related to complexity economics, given the complexity of the political economy of the euro-crisis. One recent attempt to combine agent-based macroeconomic modeling and policy analysis in Europe can be found in [50].
We also suggest to develop new models for policy advice in areas which are discussed in the general public and in politics, but currently are non-existent or extremely underrepresented in academic economics. Examples include debates about Post-/De-Growth, basic income or gender relations. The latter would require a project we propose to call “Gender-Aware Economics”, and is discussed e.g. at the Young Scholars Initiative of INET ([53]), amoung other places. Although there is a community of “feminist economics” (see e.g. [54]), it is virtually absent in most mainstream economics departments or let alone macroeconomic models used for policy advice. There has also not yet been an attempt to model the economic impact of traditional and progressive male role models (see e.g. in [64]) [^1] On p. 71 of [64], it is e.g. mentioned that the life expectancy of a low-income male member of the (German) society is 15 years lower than of a high-income women. Probably there is a bunch of very different reasons for this fact - the question is: Why is there no economic model for that, which gives a policy advice how to change it?.
@@
6 Outlook II: Progressive Data Science
In the EU, a “Circular economy Action Plan” [74] was adopted, and one central part of it is to improve the availibility of digital product data. The “EU digital product passport” (see e.g. [73] or [75]) is designed to provide information about each product’s origin, materials, environmental impact, and disposal recommendations. So in the future, a lot of (ecological) data will become available, and we propose to use it for ecological ecomomic planning. This creates the need for what we propose to call "Progressive Data Science", i.e to combine methods from data science (and AI) for a progressive change of the economic system.
@@ 6 Outlook II: Progressive Data Science
A starting point for such a project is to look at the European economy from an overall, ecological perspective. We propose to evaluate in how far new technologies and AI-methods allow for a new organization of an ecological economy at a European level: Can we move beyond the traditional way of doing business that all private companies plan only for themselves and achieve a more comprehensive ecological economic planning at the level of more aggregated entities? Recent developments in Data Science and AI (e.g. in the field of reinforcement learning) have enabled economic planning at large scales e.g. in multinational companies in a way that was unthinkable a decade ago, as it was published in the book "The people’s republic of Walmart". We propose to use such state of the art methods from AI, reinforcement learning and game tree search to evaluate the material flows in the European Economy, and to discuss the potential for ecological economic planning on a European level on that basis. This is also directly linked to plans of the EU to introduce a „Digital Product Passport“. With this European- wide product data infrastructure for all products sold and used in the EU, new forms of ecological economic planning will be become possible. The discussion about the requirements and design of digital product passes has just started in recent years. We propose to contribute to this discussion by defining requirements for a digital product passport to enable European ecological economic planning.
@@
Recently, there has been an application of data from "Life cycle sustainability assessment" to planning purposes. A starting point for the project of "Progressive Data Science" could e.g. be openLCA (see openlca.org), a free and open source software for modelling the life cycle of products and sustainability. For more context on LCA, see the recent work on "Life cycle sustainability assessment for sustainable bioeconomy, societal-ecological transformation and beyond" [76]. Ultimately, we believe a combination of Critical Mathematical Economics and Progressive Data Science is necessary in order to achieve economic democracy within planetary boundaries, as outlined in the final section of [77].

### v3

@@ Abstract
From our point of view, mathematics has been partly misused in mainstream economics to justify ‘unregulated markets’. We identify two key parts of CME, which leads to a natural structure of this article: The first part focusses on an analysis and critique of mathematical models used in mainstream economics, like e.g. the Dynamic Stochastic General Equilibrium (DSGE) in Macroeconomics and the so-called “Sonnenschein-Mantel-Debreu”-Theorems.
The aim of the second part is to improve and extend heterodox models using ingredients from modern mathematics and computer science, a method with strong relation to Complexity Economics. We exemplify this idea by describing how methods from Non-Linear Dynamics have been used in Post-Keynesian Macroeconomics’, and also discuss (Pseudo-) Goodwin cycles and possible Micro- and Mesofoundations.
Finally, we outline in which areas a collaboration between mathematicians and heterodox economists could be most promising, and discuss both existing projects in such a direction as well as areas where new models for policy advice are most needed. In an outlook, we discuss the role of (ecological) data, and the need for what we call Progressive Data Science.
@@ 1 Introduction
Also, there has been a “Workshop on Complexity and Economics” ([57]) at the Fields Institute, which was organized in collaboration with the “Young Scholars Initiative” (YSI) of INET ([55]) and especially the “YSI Complexity Economics Working Group” ([56]. The aim of this article is to describe some of the research is this direction - in order to see what mathematics and mathematicians can contribute to the development of a competitive alternative to the neoclassical paradigm in economics.
2 The Use of Mathematics in Standard Economics
@@ 3.3 A probabilistic Micro-/Meso-foundation of Macroeconomics
3.4 Collaboration between Mathematicians and Heterodox Economists
@@
From a Dynamical Systems point of view, other promising topics to look at include Hysteresis, e.g. with relation to Keynesian modelling of the labour market ([49]). In the future maybe even advanced topics like Spatio-Temporal Patterns, Turing instability or Delayed Feedback Control could be part of the mathematics used for new economic thinking - for a panorama of advanced methods in dynamical systems theory and their interplay with a wide range of applications see e.g. ([58]).
4 Conclusion: Foundations of Controversies in Economic Policy
@@
We also suggest to develop new models for policy advice in areas which are discussed in the general public and in politics, but currently are non-existent or extremely underrepresented in academic economics. Examples include debates about Post-/De-Growth, basic income or gender relations. The latter would require a project we propose to call “Gender-Aware Economics”, and is discussed e.g. at the Young Scholars Initiative of INET ([53]), amoung other places. Although there is a community of “feminist economics” (see e.g. [54]), it is virtually absent in most mainstream economics departments or let alone macroeconomic models used for policy advice. There has also not yet been an attempt to model the economic impact of traditional and progressive male role models (see e.g. in [61]).
@@
5 Outlook: Progressive Data Science
We are deeply convinced that the ecological crisis necessarily demands research on post-capitalist economic systems, to which the ’STEM subjects’ science, technology, engineering, mathematics (as well as many other fields) urgently need to contribute in order to achieve social progress. So in addition to critical mathematical economics, we propose to do research in what we call ’progressive data science’, with strong connections to (progressive) computer science and artificial intelligence. In this last section of this paper, we elaborate a bit more on this promising research area.
@@ 5 Outlook: Progressive Data Science
The aim for such an ambitious project is e.g. to look at the European economy from an overall, ecological perspective. We propose to evaluate in how far new technologies and AI-methods allow for a new organization of an ecological economy at a European level: Can we move beyond the traditional way of doing business that all private companies plan only for themselves and achieve a more comprehensive ecological economic planning at the level of more aggregated entities? Recent developments in Data Science and AI (e.g. in the field of reinforcement learning) have enabled economic planning at large scales e.g. in multinational companies in a way that was unthinkable a decade ago, as it was published in the book "The people’s republic of Walmart" [73]. We propose to use such state of the art methods from AI, reinforcement learning and game tree search to evaluate the material flows in the European Economy, and to discuss the potential for ecological economic planning on a European level on that basis. In the EU, a “Circular economy Action Plan” [71] was adopted, and one central part of it is to improve the availibility of digital product data. The “EU digital product passport” (see e.g. [70] or [72]) is designed to provide information about each product’s origin, materials, environmental impact, and disposal recommendations. With this European-wide product data infrastructure for all products sold and used in the EU, new forms of ecological economic planning will be become possible. The discussion about the requirements and design of digital product passes has just started in recent years, and we propose to define requirements for a digital product passport to enable European ecological economic planning.
So in the future, a lot of (ecological) data will become available, including ecological data e.g. for so-called “lifecycle assessment”. This creates the need for "Progressive Data Science", i.e to use methods from data science (and AI) for a progressive change of the economic system. Since a few years, there is a hype around „Artificial Intelligence“ (AI), and we propose to have a closer look on the role that digital product data and AI (in a broad definition, see below) can play for a democratic economic system that respects planetary boundaries. One focus lies naturally on on ecological questions, such as those arising in order to „close the loops“ of material flows and achieving a „circular economy“, while also the role of care work and “male role models” is very important and should be considered from the start (see [74] or [78]). The current hype of AI was fuelled by the introduction of large language models (“LLMs”) like ChatGPT, and nowadays “AI” and “LLMs” are often used interchangeably. However, for our purpose, AI can and should be defined much more broadly: A classic definition by Marvin Minsky from the 1960s defined AI as “the science of making machines do things that would require intelligence if done by humans”, and in a recent article on artificial intelligence and modern planned economies [76], Spyros Samothrakis noted that historic debates and proposals on "Computerised Economic Planning" (CEP, also see [75]) “have been inspired by and/or have touched upon other numerate disciplines like cybernetics, game theory, optimisation, complex systems, machine learning, and statistics. Arguably, the applied cutting edge of the fields that have partially contributed to the debate is currently being studied under the broad umbrella of artificial intelligence (AI)".
@@
Recently, there has been an application of data from "Life cycle sustainability assessment" to planning purposes. A starting point for the project of "Progressive Data Science" could e.g. be openLCA (see openlca.org), a free and open source software for modelling the life cycle of products and sustainability. For more context on LCA, see e.g. [77]. Ultimately, we believe a combination of Critical Mathematical Economics and Progressive Data Science is necessary in order to achieve economic democracy within planetary boundaries, as outlined e.g. in [79] or in the final section of [80].